Factor:
4x(x-1)=15
So far I have 4x^2-4x-15=0
But when it comes to factoring i"m stuck. Is this correct so far?

Answer by richwmiller(17219)

So far so good. Now you need to learn how to factor.
multiply the first coefficient 4 (of 4x^2) times the constant (-15)
4(-15)=-60
We want factors of -60 that add up to the x coefficient (-4)
-60 1
-30 2
-15 4
-10 6
wait up -10 + 6 equal -4
4x^2-4x-15=0
separate -4x into two parts (Remember FOIL)
4x^2+6x-10x-15=0
It doesn't matter which order we put (-10x+6x) or (6x-10x)
2x*(2x+3)-5*(2x+3) remember - times + is -
(2x-5)*(2x+3)=0
FOIL
4x^2-10x+6x-15
4x^2-4x-15